# Mechanically Interlocked Polymers in Dilute Solution under Shear and Extensional Flows: A Brownian Dynamics Study

### v2

@@ VII Supplementary Information
The supplementary information includes linear and MIP equilibrium properties (diffusion coefficient, relaxation time, radius of gyration, asphericity, prolateness); additional shear flow measurements as a function of Wi (gradient and cross-flow extent, orientation angle); comparisons of linear and MIPs with similar contour length; trajectory snapshots; probability distribution of rings in polyrotaxanes; and power-law exponents ($\beta$) of $\eta$ and $\psi_{1}$ fitted from shear data.
@@ Appendix B Additional Shear Flow Properties
Figure B.1: Polyrotaxane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y_{bb}\rangle/L)$, (b) fractional extension in cross direction $(\langle Z_{bb}\rangle/L)$, (c) orientation angle ($\theta$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Polyrotaxanes with 40, 80, and 120 backbone beads ($N_{bb}$) are shown in blue, green, and red, respectively; darker shades indicate higher ring density ($N_{\mathrm{rings}}/N_{bb}$). The rings in the polyrotaxanes are composed of 8 beads, regardless of backbone length.
To calculate tumbling time, we monitor the angle between end-to-end vector of the polymers with the x-axis as it progresses through the four quadrants of a full $360^{\circ}$ rotation cycle and record the time taken for the chains to complete each full cycle (i.e., one complete rotation through all four quadrants) throughout the simulation and use the mean value as the tumbling time. We use the radius of gyration along the flow and gradient directions to calculate the orientation angle $\theta$ to measure the alignment of the chains in the $xy$ plane using the following equation Schroeder et al. (2005).
@@
Figure B.2: Daisy chain properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y_{bb}\rangle/L)$, (b) fractional extension in cross direction $(\langle Z_{bb}\rangle/L)$, and (c) orientation angle ($\theta$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Daisy chains with 8, 16, and 32 beads per segment (BPS) are shown in blue, green, and red, respectively; darker shades indicate more segments ($N_{\mathrm{seg}}$). The rings in all daisy chains are composed of 8 beads.
@@
Figure B.3: Polycatenane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y_{bb}\rangle/L)$, (f) fractional extension in cross direction $(\langle Z_{bb}\rangle/L)$, and (c) orientation angle ($\theta$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Colors from cold to hot indicate higher number of rings ($N_{\mathrm{rings}}$) with fixed beads per ring ($\mathrm{BPR}=8$).
According to Fig. B3c, at low Wi polymers remain unaligned. At moderate Wi, the orientation angle $\theta$ first rises then falls, with polycatenanes of 8 beads per ring exhibiting a secondary increase previously noted in other systems Dutta and Sing (2024). A terminal decrease, reflecting alignment with very strong flow, could not be verified here because chains broke at high Wi. Having more or smaller rings shifts the primary peak to higher Wi and brings the secondary rise to lower Wi. Having more or larger rings make the orientation angle approach that of a linear polymer.
Figure B.4: Polycatenane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y_{bb}\rangle/L)$, (b) fractional extension in cross direction $(\langle Z_{bb}\rangle/L)$, and (c) orientation angle ($\theta$). Darker shades indicate higher beads per ring (BPR) in polycatenanes with fixed number of rings ($N_{\mathrm{rings}}=10)$.

### v3

@@ VII Supplementary Information
The supplementary information includes linear and MIP equilibrium properties (diffusion coefficient, relaxation time, radius of gyration, asphericity, prolateness); additional shear flow measurements as a function of Wi (gradient direction and vorticity direction extent, orientation angle); comparisons of linear and MIPs with similar contour length; comparisons of linear and MIPs with similar molecular weight with and without HI;trajectory snapshots; probability distribution of rings in polyrotaxanes; and power-law scaling exponents ($\beta$) of $\eta$, $\psi_{1}$, and $\omega\tau_{R}$ fitted from shear data.
@@ Appendix B Additional Shear Flow Properties
Figure B.1: Polyrotaxane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y\rangle/L)$, (b) fractional extension in vorticity direction $(\langle Z\rangle/L)$, (c) orientation angle ($\theta$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Polyrotaxanes with 40, 80, and 120 backbone beads ($N_{bb}$) are shown in blue, green, and red, respectively; darker shades indicate higher ring density ($N_{\mathrm{rings}}/N_{bb}$). The rings in the polyrotaxanes are composed of 8 beads, regardless of backbone length.
To calculate tumbling time, we monitor the angle between end-to-end vector of the polymers with the x-axis as it progresses through the four quadrants of a full $360^{\circ}$ rotation cycle and record the time taken for the chains to complete each full cycle (i.e., one complete rotation through all four quadrants) throughout the simulation and use the mean value as the tumbling time. We use the radius of gyration along the flow and gradient directions to calculate the orientation angle $\theta$ to measure the alignment of the chains in the $xy$ plane using the following equation schroeder2005dynamics.
@@
Figure B.2: Daisy chain properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y\rangle/L)$, (b) fractional extension in vorticity direction $(\langle Z\rangle/L)$, and (c) orientation angle ($\theta$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Daisy chains with 8, 16, and 32 beads per segment (BPS) are shown in blue, green, and red, respectively; darker shades indicate more segments ($N_{\mathrm{seg}}$). The rings in all daisy chains are composed of 8 beads.
@@
Figure B.3: Polycatenane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y\rangle/L)$, (f) fractional extension in vorticity direction $(\langle Z\rangle/L)$, and (c) orientation angle ($\theta$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Colors from cold to hot indicate higher number of rings ($N_{\mathrm{rings}}$) with fixed beads per ring ($\mathrm{BPR}=8$).
According to Fig. B3c, at low Wi polymers remain unaligned. At moderate Wi, the orientation angle $\theta$ first rises then falls, with polycatenanes of 8 beads per ring exhibiting a secondary increase previously noted in other systems dutta2024brownian. A terminal decrease, reflecting alignment with very strong flow, could not be verified here because chains broke at high Wi. Having more or smaller rings shifts the primary peak to higher Wi and brings the secondary rise to lower Wi. Having more or larger rings make the orientation angle approach that of a linear polymer.
Figure B.4: Polycatenane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y\rangle/L)$, (b) fractional extension in vorticity direction $(\langle Z\rangle/L)$, and (c) orientation angle ($\theta$). Darker shades indicate higher beads per ring (BPR) in polycatenanes with fixed number of rings ($N_{\mathrm{rings}}=10)$.
@@
Figure B.6: Comparison of MIPs of similar molecular weight in extensional (a, b) and shear (c, d) flow. All properties are normalized by linear polymer (solid horizontal line). Fractional extension in the flow direction $(\langle X_{bb}\rangle/L)$ for extension (a) and shear (c) along with extensional viscosity $(\eta_{e})$ (b) and shear viscosity $(\eta)$ are given as a function of Weissenberg number (Wi). The linear polymer (black line) has 80 backbone beads ($M_{w}=80$). The polyrotaxane (blue) has 40 backbone beads including 2 capping beads and 6 rings ($M_{w}=88$). The daisy chain (red) has 4 linear segments of 8 beads each interlocked with 6 8 bead rings ($M_{w}=74$) and includes 2 capping beads. The polycatenane (green) has 10 interconnected rings of 8 beads each ($M_{w}=80$).
Figure B.7: Comparison of MIPs of similar molecular weight in extensional (a, b) and shear (c, d) flow with (solid lines and filled markers) and without hydrodynamic interaction (dashed lines and hollow markers). Fractional extension in the flow direction $(\langle X_{bb}\rangle/L)$ for extension (a) and shear (c) along with extensional viscosity $(\eta_{e})$ (b) and shear viscosity $(\eta)$ are given as a function of Weissenberg number (Wi). The linear polymer (black line) has 80 backbone beads ($M_{w}=80$). The polyrotaxane (blue) has 40 backbone beads including 2 capping beads and 6 rings ($M_{w}=88$). The daisy chain (red) has 4 linear segments of 8 beads each interlocked with 6 8 bead rings ($M_{w}=74$) and includes 2 capping beads. The polycatenane (green) has 10 interconnected rings of 8 beads each ($M_{w}=80$).
@@ Appendix E Power-law exponents
Table E.3: Power-law exponents ($\beta$) of $\omega\tau_{R}$ for different polymer architectures
Architecture $M_{w}$ $N_{bb}$ BPS BPR $N_{\mathrm{rings}}$ $N_{\mathrm{seg}}$ $\beta$
Linear 40 - - - - - 0.741 $\pm$ 0.008
Linear 80 - - - - - 0.752 $\pm$ 0.016
Linear 120 - - - - - 0.732 $\pm$ 0.006
Polyrotaxane 56 40 - - 2 - 0.740 $\pm$ 0.011
Polyrotaxane 72 40 - - 4 - 0.751 $\pm$ 0.014
Polyrotaxane 104 40 - - 8 - 0.826 $\pm$ 0.011
Polyrotaxane 112 80 - - 4 - 0.748 $\pm$ 0.020
Polyrotaxane 144 80 - - 8 - 0.763 $\pm$ 0.017
Polyrotaxane 208 80 - - 16 - 0.794 $\pm$ 0.012
Polyrotaxane 168 120 - - 6 - 0.766 $\pm$ 0.013
Polyrotaxane 216 120 - - 12 - 0.770 $\pm$ 0.015
Polyrotaxane 312 120 - - 24 - 0.787 $\pm$ 0.018
Daisy chain 30 - 8 - - 2 0.838 $\pm$ 0.006
Daisy chain 74 - 8 - - 4 0.822 $\pm$ 0.009
Daisy chain 118 - 8 - - 6 0.838 $\pm$ 0.011
Daisy chain 206 - 8 - - 10 0.867 $\pm$ 0.016
Daisy chain 46 - 16 - - 2 0.875 $\pm$ 0.021
Daisy chain 106 - 16 - - 4 0.864 $\pm$ 0.018
Daisy chain 226 - 16 - - 8 0.882 $\pm$ 0.022
Daisy chain 78 - 32 - - 2 0.834 $\pm$ 0.022
Daisy chain 170 - 32 - - 4 0.822 $\pm$ 0.021
Daisy chain 262 - 32 - - 6 0.840 $\pm$ 0.021
Polycatenane 40 - - 8 5 - 0.917 $\pm$ 0.007
Polycatenane 80 - - 8 10 - 0.935 $\pm$ 0.010
Polycatenane 168 - - 8 21 - 0.912 $\pm$ 0.030
Polycatenane 344 - - 8 43 - 0.989 $\pm$ 0.013
Polycatenane 520 - - 8 65 - 1.029 $\pm$ 0.015
Polycatenane 160 - - 16 10 - 0.790 $\pm$ 0.029
Polycatenane 240 - - 24 10 - 0.747 $\pm$ 0.031
Polycatenane 320 - - 32 10 - 0.753 $\pm$ 0.021